Thanks so much for all your help . After see the problem solve i try to work on them to make sure the are right . And i try to see hoe you all get the answer it is helping me to understand my probles .
So if some one can help me with these problems .
1) solve for x .collect xs on the LEFT side .The solution be presented with x first the correct inequality symbol ,then the number .
I have 10 of these to do .
but please help me with these .
1) 5(2x -1) <-8
2) X + 4 -8x-1≥3(x-2)


3) Name the coordinate described here
From the origin, the point is three places to the right and eight places down..

4) Find the midpoint between these points:
(-2, 7) and (-8,-5)


5) Please help me fill in the chart given this linear equation
Y=5x-4
I have to type all work below in the chart .Please I have 5 of these to do just give me an example on this one more thing where can I get the chart.



6) The linear equation y = 6x-5 graphs as a horizontal/ vertical/ diagonal line (choose the correct label and type it below)



Please tell me where I can get to do the midpoint the art for it.
Thanks .Enida

Answer by stanbon(75887)

1) 5(2x -1) <-8
2x-1 <= -8/5
Add 1 or 5/5 to both sides to get:
2x <= -3/5
Divide both sides by 2 to get:
x <= -3/10
--------------
2) X + 4 -8x-1≥3(x-2)-
-7x + 3 >= 3x-6
Add 7x to both sides
3 > = 10x-6
Add 6 to both sides to get:
9 >=10x
Divide both sides by 10 and
Reverse the statement to get:
x<= 9/10
----------------

3) Name the coordinate described here
From the origin, the point is three places to the right and eight places down.
(3,-8)
-----------.
4) Find the midpoint between these points:
(-2, 7) and (-8,-5)
x coordinate : (-2-8)/2 = -5
y coordinate : (7-5)/2 = 1
(-5,1) is the midpoint
-----------------------
5) Please help me fill in the chart given this linear equation
Y=5x-4
I have to type all work below in the chart .Please I have 5 of these to do just give me an example on this one more thing where can I get the chart.
If x=0, y=-4
If x=1, y=1
If x=2, y=6
etc.
------------

6) The linear equation y = 6x-5 graphs as a horizontal/ vertical/ diagonal line (choose the correct label and type it below)
diagonal because the slope is +6
-----------
Cheers,
Stan H.